# MapleGrasp: Mask-guided Feature Pooling for Language-driven Efficient Robotic Grasping

### v1

@@ 4 RefGraspNet Dataset: Learning to Grasp from 200M+ Grasp Poses
Table 1: Comparison of existing 4-DoF LDRG datasets. RefGraspNet provides 200M+ high quality grasps, object masks for each scene, and thus can be used to finetune both referring segmentation and grasping models.
We build on existing research and datasets in language driven grasping by creating a new corpus of real world cluttered images and their grasp annotations. Our dataset, RefGraspNet, is derived from the GraspNet-1B dataset [11] and comprises a diverse corpus collected using 88 unique objects arranged in cluttered scenes. We construct 10 instruction templates such as “Grasp the {object}”, “Find the {object} in this scene”, etc. and expand the object names to create referring expressions. In scenes with multiple objects, we leverage an open-source model, DeepSeek-VL [30], to provide additional attributes such as color, shape and relative position. We chose GraspNet-1B as our base corpus since it consists of 1 billion + grasp poses annotated with sampling and force-closure techniques. We sample grasps with a confidence score greater than 70%, and our final corpus comprises 219M grasp poses, 8 times bigger than the previous largest dataset for language driven 4 DoF grasping. Comparisons with prior datasets are provided in Tab. 1. We randomly choose 70% of the objects to be seen objects (with the rest as unseen) and create train/val/test-seen and test-unseen splits for our experiments.
@@ 5 Experiments
Here, we present our experiments evaluating MapleGrasp on OCID-VLG and RefGraspNet benchmarks, while also comparing our framework to prior LDRG and VLA methods in simulation and real-world setups.
5.1 OCID-VLG Benchmark
The OCID-VLG corpus [46] contains 1763 highly cluttered indoor tabletop RGB-D scenes with 31 unique objects. Objects have annotated 4-DoF grasp poses and segmentation masks, associated with referring expressions describing attributes such as color, shape, and relative position. With 89K unique (RGB-Text-Mask) tuples, it is the largest publicly available benchmark for language-driven grasp detection frameworks [^1] GraspAnything++ is another open-source corpus, but the publicly shared subset is significantly smaller than what was used for training their models, preventing fair benchmarking.. We evaluate various CLIP-based architectures on OCID-VLG, applying the MapleGrasp framework to Ref2Grab [2], which uses frozen CLIP encoders for improved generalization, and CROG [46], as detailed in Sec. 3.4. Results are summarized in Tab. 2.
Model IOU Pr@50 Pr@60 Pr@70 Pr@80 Pr90 J@1 J@Any
DetSeg + CLIP 29 27.2 20.9 17.5 17.2 16 28.1 39.2
GR-ConvNet + CLIP 31.3 21.0 11.6 5.5 2.4 0.5 9.73 15.41
SSG + CLIP 33.6 35.6 35.6 35.5 35.5 32.8 33.5 34.7
CROG 81.1 96.9 94.8 87.2 64.1 16.4 77.2 87.7
Ref2Grab 83.26 85.68 85.13 91.53 89.69 83.21 70.5 79.1
ETRG 80.11 - - - - - 82.28 91.12
MapleGrasp-Ref2Grab 83.78 85.96 85.41 91.74 90.13 83.79 76.8 84.7
MapleGrasp-CROG 81.36 97.40 95.32 87.9 65.4 16.4 86.15 91.9
Table 2: Results on OCID-VLG. We implement all baseline methods except ETRG, whose results are taken from their paper. Our comparison is limited to methods that do not utilize depth maps during training.
@@ 5.1 OCID-VLG Benchmark
Table 3: Qualitative comparisons of LDRG approaches under complex referring queries. The zero-shot baseline (first column) struggles to identify referred object due to distractors. MapleGrasp (last column) can predict a stable and accessible grasp for each queried object.
Applying mask proposals enhances both the CROG and Ref2Grab baselines in terms of the Jacquard Index J@N metric, which evaluates the top-N grasp rectangle candidates that exhibit a rotation angle difference of less than $30^{\circ}$ and an IoU greater than 0.25 compared to the ground-truth grasp rectangle. MapleGrasp achieves a 12% relative improvement in J@1 over previous state-of-the-art methods, demonstrating higher precision in grasp prediction. However, referring expression segmentation (RES) accuracy, measured using Prec@X, which quantifies the percentage of test cases where the IoU exceeds $X\in\{0.5,0.6,0.7,0.8,0.9\}$, remains relatively unchanged. This may be due to CROG losing pretrained knowledge during CLIP layer fine-tuning. In contrast, MapleGrasp-Ref2Grab, which employs frozen vision-language encoders, achieves higher performance in the RES task but struggles with grasp map prediction due to its limited trainable parameter space (4M vs. CROG’s 167M). Overall, our findings suggest that MFP enhances both frozen and fully fine-tuned CLIP-based models for language-driven grasp detection.
5.2 Ablation Studies
@@
We conduct ablation studies on MapleGrasp (LABEL:tab:ablations) to analyze its effects due to architectural modifications. Results highlight the importance of accurate RES for mask pooling; using Molmo, which achieves lower IoU on RES tasks, negatively impacts grasp accuracy due to complex referring queries requiring task-specific visual grounding. While using frozen CLIP improves RES, it decreases grasp performance because of fewer trainable parameters. Replacing cross-attention with MLP-Mixers, despite reported benefits [44], does not significantly improve performance. Furthermore, replacing the weighted smooth L1 loss with a standard Smooth L1 loss leads to reduced scores, likely from slower convergence. All ablations use a fixed number of training epochs for consistency.
Method IoU J@1 J@Any
MapleGrasp-CROG 81.36 86.15 91.90
(1) Molmo+SAM for Mask Pooling 73.42 77.79 80.16
(2) Frozen CLIP Layers 83.52 72.19 74.44
(3) No Cross-Attn, MLP-Mixer 80.16 84.37 90.45
(4) Standard Smooth L1 Loss 80.17 83.23 89.18
5.3 RefGraspNet Benchmark
We further evaluate previous baselines and MFP-enhanced architectures on our RefGraspNet dataset, establishing a strong benchmark for future research. The dataset comprises 88 unique objects, with 70% assigned as seen objects and 30% as unseen objects. Results for both test splits are reported in Tab. 4. Similar trends as observed in the OCID-VLG benchmark hold here, with MFP providing significant performance gains for both Ref2Grab and CROG. However, all trained models perform worse on unseen objects, highlighting a lack of zero-shot generalization. In contrast, open-set object detectors such as Molmo [7] achieve higher test unseen IoU. This motivates us to incorporate Molmo in an ensembling setup to boost zero-shot segmentation, which is essential for robust grasp pose estimation.
Test Seen Test Unseen
Model IoU J@1 J@Any IoU J@1 J@Any
DetSeg + CLIP 48.18 40.19 41.56 30.57 27.18 27.20
GR-ConvNet +
CLIP 52.34 44.19 46.56 32.18 30.19 32.17
SSG + CLIP 57.16 48.78 50.10 38.19 34.16 34.89
CROG 84.49 85.32 86.49 69.95 70.81 71.99
Ref2Grab 87.32 73.27 75.45 75.67 62.12 63.28
MapleGrasp-Ref2Grab 87.74 75.67 77.89 75.11 65.19 67.22
MapleGrasp-CROG 84.86 89.15 89.67 70.43 74.92 75.67
Table 4: Results on RefGraspNet. The MFP framework leads to consistent improves across seen and unseen objects. However, performance on unseen objects deteriorates for all methods.
@@ 5.4 Cross-Dataset Testing
To assess the generalizability of models trained on RefGraspNet, we cross-test them on the OCID-VLG benchmark. We compare MapleGrasp-Ref2Grab and MapleGrasp-CROG, and present results in Tab. 5. Our findings reveal that models trained on RefGraspNet generalize better than those trained solely on OCID-VLG. The dataset’s diversity in grasp poses and text expressions enhances inference on other benchmarks. This highlights the advantages of training on RefGraspNet, encouraging further adoption of our corpus to push performance boundaries.
@@
Model Train-Dataset J@1 J@Any J@1 J@Any
MapleGrasp-Ref2Gr OCID-VLG 76.8 84.7 41.8 42.9
RefGraspNet 71.2 74.4 75.7 77.9
MapleGrasp-CROG OCID-VLG 86.2 91.9 33.2 35.6
RefGraspNet 68.2 73.8 89.2 89.7
Table 5: Cross-dataset evaluation results. Models trained on RefGraspNet generalize better than OCID-VLG.
5.5 LIBERO Simulation Experiments
We deploy the MapleGrasp model trained on RefGraspNet and extend grasp poses to 3D using an inverse kinematics controller for trajectory prediction (Sec. 3.6). This setup allows direct comparisons with Vision-Language-Action (VLA) models. Since our approach focuses on language-driven grasp pose estimation, we cannot evaluate task success rates directly. Instead, we define grasp success as trials in which an object is lifted and held above ground level for at least two seconds. We report J@1 and J@Any metrics. Experiments are conducted on three task suites from the LIBERO benchmark: SPATIAL, GOAL, and OBJECT. Since VLAs are fine-tuned on 50 teleoperated demonstrations per task, we automate 4-DoF grasp annotations using forward kinematics-based 3D grasp pose estimation and project the masks to 2D, inverting steps described in Sec. 3.1. We fine-tuned MapleGrasp on these annotations to adapt to the simulation environment. At inference time, MapleGrasp predicts 2D grasps, which are transformed and executed in 3D by the controller. Results in Tab. 6 compare our method with VLA and diffusion baselines.
@@ 5.5 LIBERO Simulation Experiments
Diffusion Policy 0.77 0.69 0.92
Octo 0.78 0.84 0.88
OpenVLA 0.84 0.79 0.90
MapleGrasp-CROG 0.87 0.85 0.90
Table 6: Libero simulation results across SPATIAL, GOAL and OBJECT task suites. Scores are reported using J@1.
Our results demonstrate that our grasping+control pipeline achieves similar accuracy to recent VLAs in grasping tasks. Interestingly, decoupling grasping and control improves generalization, as trajectory-based imitation learning methods fail to complete unseen tasks. However, our model successfully determines grasp poses for similar but unseen objects, allowing the control algorithm to generate a stable trajectory without requiring video demonstrations. Cross-dataset testing in Tab. 7 reveals the capabilities of our approach in handling unseen queries across task suites. Since our LDRG model is only concerned with predicting grasp poses and we leverage a controller to determine the trajectory, our method can be deployed without learning from tele-operated demonstrations.
@@
OpenVLA GOAL 0.79 0.00
OBJECT 0.00 0.90
MapleGrasp-CROG GOAL 0.85 0.62
OBJECT 0.68 0.90
Table 7: Comparison of OpenVLA and our method in cross-dataset settings. OpenVLA struggles with seen objects under novel tasks, while our method generalizes well to unseen queries.
@@ 5.6 Real-Robot Experiments
Figure 3: Setup for robot trials using 7 DoF franka arm and our modular pipeline described in section Sec. 3.6
We evaluate various LDRG methods in a real-world robotic setup using a Franka arm with a mounted camera at the gripper, as shown in Fig. 3. We construct diverse tabletop scenes featuring single and multiple objects, distractors, and varying lighting conditions. Results are presented in Tab. 8. For real-robot trials, MapleGrasp integrates Molmo-based object mask refinement, which compares our model’s segmentation predictions with Molmo’s out-of-vocabulary predictions, selecting the highest mutual overlap. MapleGrasp outperforms other baselines, achieving a notable 7% relative improvement in cluttered setups.
Model Single Cluttered
Det-Seg+CLIP 0.27 0.21
GR-ConvNet+CLIP 0.30 0.27
Molmo+SAM+GraspNet 0.45 0.47
SSG+CLIP 0.21 0.17
Ref2Grab 0.42 0.40
CROG 0.55 0.53
MapleGrasp-CROG 0.59 0.57

### v2

@@ 3.3 MapleGrasp
Table 1: Comparison of existing LDRG datasets. RefGraspNet provides 200M+ high quality grasps, object masks for each scene, and thus can be used to finetune both referring segmentation and grasping models.
3.4 Loss Function
Prior work utilizes the Smooth L1 loss for pixel-wise prediction in visual grounding, defined per pixel $(\hat{x}_{p},x_{p})$ as:
Here, $\hat{x}_{p}\in\mathbb{R}$ denotes predicted value at pixel $p$, $x_{p}\in\mathbb{R}$ is the ground-truth value, and $\beta>0$ is a constant that determines the transition point between L2 and L1 loss regions.
3.5 Metrics
Language-driven grasp poses are evaluated based on their correctness, feasibility, and overall quality, given a language instruction and the corresponding scene image. Quantitative evaluation is performed using grasp success rate for both Top-1 and Top-K predictions. For 4-DoF grasping, a prediction is considered successful if its rotation angle differs from the ground-truth by less than $30^{\circ}$, and its intersection-over-union (IoU) with the ground-truth grasp rectangle exceeds 0.25 [31, 45]. For 6-DoF grasping, success is determined if the selected grasp proposal corresponds to the ground-truth grasp for the referred object [8].
4 RefGraspNet Dataset: Learning to Grasp from 200M+ Grasp Poses
To advance language-driven robotic grasping, we introduce RefGraspNet, a richly annotated dataset of over 219 million high-quality 6 DoF grasp poses, derived from real-world cluttered scenes. Building on GraspNet-1B [10], which provides over one billion sampled grasps with force-closure scoring comprising 88 distinct household and industrial objects arranged in hundreds of densely cluttered table-top scenarios. Each scene captured from multiple calibrated viewpoints yielded object masks and 6 DoF grasp poses.
We automate generating referring expressions using DeepSeek-VL [28], an open-source VLM. Starting from ten instruction templates (e.g., “Grasp the {object},” “Locate the {object}”), we expanded object mentions with color, shape, and spatial qualifiers in multi-object layouts (e.g., “Grasp the red cup behind the yellow bowl”). This yielded over 12.25M unique referring phrases, ensuring robust grounding of grasp commands in cluttered environments. Grasp candidates were filtered using a 70% force-closure confidence threshold, which was selected based on extensive manual validation, resulting in 219 million validated 6 DoF grasps—roughly eight times larger than the previous largest language-driven grasp dataset. We provide comprehensive comparisons in Tab. 1, demonstrating improvements in scale, scene complexity, and linguistic variety. To support empirical studies, we split our objects into 70% “seen” (62 objects) and 30% “unseen” (26 objects), and generate train/val, test-seen, and test-unseen splits. This design enables evaluation of generalization to novel objects and arrangements. Examples are shown in Fig. 4.
Figure 4: Samples from RefGraspNet showcasing cluttered scenes with multiple viewpoints, with diverse objects, distractors, and grasping instructions. Only top-down grasps shown for brevity.
Model Epochs to Conv. OCID-VLG RefGraspNet
Test Seen Test Unseen
DetSeg + CLIP [1] 22 28.12 / 39.21 40.19 / 41.56 27.18 / 27.20
GR-ConvNet + CLIP [20] 15 9.73 / 15.41 34.19 / 46.56 30.19 / 32.17
SSG + CLIP [52] 41 33.51 / 34.70 48.78 / 50.10 14.36 / 14.89
Molmo+SAM2+GraspNet 63.22 / 69.14 68.85 / 70.27 66.23 / 69.41
CROG [44] 50 77.22 / 87.71 85.32 / 86.49 70.81 / 71.99
HiFi-CS [2] 44 70.54 / 79.12 73.27 / 74.55 62.13 / 63.28
ETRG [55] 49 82.28 / 91.12 86.13/88.12 68.11 / 70.05
Ours: MapleGrasp 32 88.15 / 92.98 89.86 / 91.67 76.92 / 77.67
Table 2: Benchmarking MapleGrasp across OCID-VLG and RefGraspNet datasets for 4-DoF and 6-DoF vision-language driven grasping respectively. Scores reported as grasping accuracy for top-1/top-5 predictions (see Sec. 3.5).
@@ 5 Experiments
We evaluate MapleGrasp on OCID-VLG and RefGraspNet, while comparing it to prior LDRG and VLA methods in both simulation and real-world settings.
5.1 Benchmark Comparisons
The OCID-VLG corpus [44] contains 1763 highly cluttered indoor tabletop RGB-D scenes with 31 unique objects. Objects have annotated 4-DoF grasp poses and segmentation masks, associated with referring expressions describing attributes such as color, shape, and relative position. With 89K unique (RGB-Text-Mask) tuples, it is the largest publicly available dataset for LDRG [^1] GraspAnything++ is open-sourced; however, its public subset is much smaller than the full training data, making fair benchmarking impossible.. We also provide a benchmark on our new contributed RefGraspNet corpus.
We evaluate various CLIP-based architectures on both datasets: CROG [44] utilizes CLIP embeddings with cross-attention, followed by MLP layers, to simultaneously predict object masks and grasp maps. HiFi-CS [2] employs a frozen CLIP model with FiLM-based fusion and upsampling to generate object masks, followed by grasp estimation using a separate frozen network. ETRG [55] introduces trainable adapters within the CLIP model to adapt it for grasping, and is trained on additional data to segment referring expressions and predict referring affordance. All baselines are trained for 50 epochs to ensure a fair comparison (see Tab. 2). We also include a strong zero-shot baseline that combines Molmo [6] for identifying the referred object, SAM 2 [37] for generating the object mask, and GraspNet [10] for predicting both 4-DoF and 6-DoF grasp poses. MapleGrasp achieves a 7% relative improvement in Top-1 success rate over previous state-of-the-art methods, demonstrating higher precision in grasp prediction in OCID-VLG. In RefGraspNet, MapleGrasp outperforms baselines for both seen and unseen objects, with a larger performance margin observed in the unseen setting. We attribute this to our closely coupled dual-stage training, which is able to provide partial object masks in unseen objects, which is then used by the grasp detection module to generate graspable object regions and accurate grasp poses. MapleGrasp trains more efficiently, converging in 32 epochs ($\sim$8 hours), compared to 45–50 epochs (12–16 hours) required by other baselines on an RTX 4090 GPU.
@@ 5.1 Benchmark Comparisons
Pass the tissues to the rear left of the blue and black marker [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Table 3: Qualitative comparisons of LDRG approaches under complex referring queries. The zero-shot baseline (first column) struggles to identify referred object due to distractors. MapleGrasp (last column) can predict an accurate and accessible grasp for queried objects in the first three rows. Last row shows an example where all models fail to recognize the correct object (ground truth in green).
5.2 Qualitative and ablation studies
We provide comparisons in (Tab. 3). In row 1 and 3, we observe that the off-the-shelf approach fails to accurately predict a grasp on the referred object. This occurs because the query contains attributes like “rightmost” and “furthest” increasing the complexity of referring. Since the masks are predicted incorrectly, the downstream GraspNet predicts an incorrect grasp. While CROG recognizes these objects better, having seen a variety of referring attributes during training, often fails to localize the grasp within the target object. MapleGrasp’s mask-guided predictions are accurate in first three rows, but fail for occluded objects and ambiguous queries, where the mask alone cannot localize the grasp pose (row 4). More examples in supplementary Sec. $\mathcal{S}$2.
Figure 5: Comparisons across text inputs: from level 1 with just object names to level 4 with color, shape, position attributes
To compare methods across language complexity, we group test queries by the number of object attributes they mention. For example, “Grasp the red circular box near the top right of the image” contains three attributes that an LDRG model must exploit to locate a grasp. Using a popular named entity recognition model, GLiNER [57], we count attributes in the RefGraspNet test split and assign queries to four bands—L1 EASY (one attribute) through L4 HARD (four). Fig. 5 plots Top-1 accuracy for MapleGrasp and two baselines: accuracy declines with added attributes, yet MapleGrasp consistently remains ahead.
@@ 5.2 Qualitative and ablation studies
We conduct ablation studies on MapleGrasp (Tab. 4) to analyze its effects due to architectural modifications. Results highlight the importance of accurate RES for mask pooling; using Molmo, which achieves lower IoU (50-60%) on object mask detection, negatively impacts grasp accuracy due to complex referring queries requiring task-specific visual grounding. While using frozen CLIP improves object identification, it reduces grasp scores due to insufficient grounding of VL features for grasp detection. Replacing cross-attention with MLP-Mixers [42], does not improve performance. Furthermore, replacing the weighted smooth L1 loss with a standard smooth L1 loss leads to reduced scores due to slower convergence. All ablations use a fixed number of training epochs.
Method Top-1 Top-5
MapleGrasp 88.15 92.90
(1) Molmo+SAM for Mask Pooling 78.69 81.27
(2) Frozen CLIP Layers 73.14 75.57
(3) No Cross-Attn, MLP-Mixer 85.42 91.54
(4) Standard Smooth L1 Loss 84.35 90.23
Table 4: Ablation study on the OCID-VLG corpus.
Figure 6: Training data efficiency comparisons across baselines on RefGraspNet. Scores reported as grasping success rate in %.
Lastly, we evaluate the training efficiency of MapleGrasp in comparison to two baseline methods. In this experiment, we vary the size of the training dataset and assess model performance after training on 20%, 40%, 60%, 80%, and 100% of the RefGraspNet corpus. The results, presented in Fig. 6, demonstrate that MapleGrasp consistently outperforms the baselines on the test set as the dataset size increases, with particularly notable improvements observed after training on just 20–40% of the data. These findings suggest that incorporating referring object masks for pooling vision-language feature maps in grasp detection enhances both training effectiveness and sample efficiency.
@@ 5.3 Cross-Dataset Testing
To assess generalizability of models trained on RefGraspNet, we cross-test them on the OCID-VLG benchmark. We compare CROG and MapleGrasp Tab. 5 and our findings reveal that models trained on RefGraspNet generalize better than those trained solely on OCID-VLG.
@@
Model Train-Dataset Top-1 Top-5 Top-1 Top-5
CROG OCID-VLG 77.2 87.7 41.8 42.9
RefGraspNet 68.2 73.8 85.32 86.49
MapleGrasp OCID-VLG 88.2 92.9 43.2 45.6
RefGraspNet 79.5 81.7 89.2 89.7
Table 5: Cross-dataset evaluation results (highlighted). Models trained on RefGraspNet generalize better than OCID-VLG.
5.4 LIBERO Physics Simulation Experiments
We test MapleGrasp in the LIBERO simulation environment in a table-top setting. The grasp poses generated by MapleGrasp are executed with an inverse kinematics velocity controller described in supplementary Sec. $\mathcal{S}$3. We compare our modular grasping+trajectory optimization method with end-to-end VLAs. While VLAs are trained on tele-operated demonstrations of successful tasks, we use these observations to identify stable grasp poses for referred objects. We only benchmark the object grasping skill. For example, consider the task “Put the wine bottle on the rack.” We consider the task to be successful if the robot grasps the “wine bottle” and lift it above the ground. Success rate is measured manually and averaged with 10 trial runs per task.
@@ 5.4 LIBERO Physics Simulation Experiments
Diffusion Policy 77 69 92
Octo 78 84 88
OpenVLA 84 79 90
MapleGrasp 87 85 90
Table 6: Libero simulation results across SPATIAL, GOAL and OBJECT task suites. Scores are reported using success rate
Experiments are conducted on three task suites from the LIBERO benchmark: SPATIAL, GOAL, and OBJECT. For more details about the tasks, refer to supplementary Sec.  $\mathcal{S}$4. Results in Tab. 6 compare our method with VLA and diffusion baselines. Our results demonstrate that our grasping+control pipeline achieves similar accuracy to recent VLAs in grasping tasks. Interestingly, decoupling grasping and control improves generalization, as trajectory-based imitation learning methods fail to complete unseen tasks. However, our model successfully determines grasp poses for similar but unseen objects, allowing the control algorithm to generate a stable trajectory without requiring video demonstrations. Cross-dataset testing in Tab. 7 reveals the capabilities of our approach in handling unseen queries across task suites. Since our LDRG model is only concerned with predicting grasp poses and we leverage a controller to determine the trajectory, our method can be deployed without learning from tele-operated demonstrations.
@@
OpenVLA GOAL 79 0
OBJECT 12 90
MapleGrasp GOAL 85 62
OBJECT 68 90
Table 7: Cross-dataset results show OpenVLA struggles on seen objects in new tasks, while MapleGrasp generalizes better.
@@ 5.5 Real-Robot Experiments
Figure 7: Setup for robot trials using 7 DoF franka arm.
We evaluate various LDRG methods in a real-world robotic setup using a Franka arm with a camera mounted at the gripper, as shown in Fig. 7. All models were fine-tuned on a combined dataset of RefGraspNet and OCID-VLG. For evaluation, we collected 15 objects, 10 seen and 5 unseen to create randomly arranged tabletop scenes (see supplementary Sec.  $\mathcal{S}$5). To test capabilities of all baselines, we performed separate experiments trying to grasp seen and unseen objects in the presence of distractors (objects that look similar to the target object either in shape, color, or type). Each grasping instruction was tested 10 times and averaged. This resulted in N = 300 (15x2x10) robot trials. Results are presented in Tab. 8. We consider a trial to be successful if the robot grasps object and lifts it 30cm above the surface.
[…]
Seen Objects Unseen objects
Model Single Cluttered Single Cluttered
Molmo+SAM2+GraspNet 70 61 71 66
HiFi-CS 72 63 65 59
CROG 85 73 69 54
MapleGrasp 93 87 72 65
MapleGrasp+Molmo 91 85 74 73